OpenStudy (anonymous):

IF X=18 (+ or -) 0.5 and Y= 9.0 (+ or -) 0.4, find appropriate values, including errors, for: X+Y X-Y X+2Y 2X-3Y X * Y X / Y sqr(XY XY^3

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

If we wanted to include the endpoints we would write [2,4]. If we wanted to ...... \frac{1}{xy^5}. 35. \sqrt{x^2y^4}. xy2. 36. \bigg(\frac{8x^6}{y^4}\. \frac{2x^2}{y^{4/3}}

OpenStudy (anonymous):

2.5 Mean Square Approximation and Parseval's Identity 16 .... ((x + 3)( U−3 − U−2) + (2x + 5)( U−2 − U−1) + (x + 4)( U−1 − U0) ... 18 x cos 3x + ∑. 6 n=1, n=3 sin nx n(9−n2) . # 49: yp = ... ↔ yp = x(.... # 67: y = − .... Same reasoning as in the previous exercise, we find the solution ...... So u(xy) = T if and only if 2xy = T if and only

OpenStudy (anonymous):

do u get it?

OpenStudy (anonymous):

this doesnt make any sense, are you sure your doing it the right way

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

here let me put it another way

OpenStudy (anonymous):

all we have to do is add X +Y for the first one 18 + 9.0 = 27 then do you just add the uncertainties? .5+.4= .9 27 + or - .9 ???

OpenStudy (anonymous):

If we wanted to include the endpoints we would write [2,4]. 2.5 Mean Square Approximation and Parseval's Identity 16 .... ((x + 3)( U−3 − U−2) + (2x + 5)( U−2 − U−1) + (x + 4)( U−1 − U0) ... 18 x cos 3x + ∑. 6 n=1, n=3 sin nx n(9−n2) .

OpenStudy (anonymous):

Same reasoning as in the previous exercise, we find the solution ...... So u(xy) = T if and only if 2xy = T if and only

OpenStudy (anonymous):

No this is none of that, this is physics lol

OpenStudy (anonymous):

is my answer correct?

OpenStudy (anonymous):

Of course. yes yes yes

OpenStudy (anonymous):

ok so you add the uncertainties?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

x+2Y = 36 + or - 0.9??